### v1

@@ 2.4 Implementation details
All experiments were conducted on an NVIDIA GeForce RTX 2080 Ti (12GB VRAM) using CUDA 12.2 and implemented in PyTorch neural network framework. The dataset was divided into 17 numerical phantoms for training, 8 for validation, and 20 for testing. Instead of using full-resolution images during training, a volumetric patch-based sampling strategy was adopted to optimize memory usage and enhance the model’s ability to capture localized PR effects. Each epoch included 4200 patches sampled for training and 1800 patches sampled for validation. Patches of three different sizes (24×24×24, 18×18×18, 12×12×12 voxels) were randomly extracted to promote multi-scale learning and improve generalization [46]. To prevent empty patches and improve robustness, a filtering step was applied to exclude empty regions based on the $\mu$-map. Given the physical constraints of PR effects, only random flipping in all three spatial directions (x, y, z) was applied. Scaling or zooming were avoided, as they could alter the PR effect, leading to inconsistencies in the network’s learning process.
@@
The network was optimized using the AdamW optimizer [47], with an initial learning rate of $10^{-3}$, and a OneCycleLR scheduler [48] dynamically adjusting learning rates and momentum to enhance training stability. The batch size was set to 120. Early stopping was applied with a patience of 50 epochs and a minimum required improvement of $10^{-7}$ in the validation loss.
@@
To handle large volumetric PET datasets efficiently, the HDF5 file format [49] was used for data storage and sampling, allowing efficient random patch extraction without loading all the patches into memory, faster data access, and reduced I/O bottlenecks during training, as well as optimized GPU memory usage [50], enabling training on hardware with limited VRAM. Unlike some implementations that use mixed precision training to reduce memory usage, our approach was trained using full precision (FP32) to ensure numerical stability, preventing potential precision errors.
@@
In this work, we leverage the properties of convolutional layers and their receptive fields to infer full images without patching [51]. Convolutional operations learn spatial patterns locally and apply them uniformly across larger images due to shared weights and growing receptive fields. Additionally, spatial equivariance ensures that features learned in one region remain valid elsewhere [52]. By integrating volumetric patches during training and inferring full images during testing, our approach enhances efficiency while maintaining consistent feature extraction across the entire volume.
@@ 2.5 Validation
Validation was performed using simulated PET datasets, where both 68Ga and 18F PET images were generated under same conditions. The network was trained to predict the reconstructed 18F-equivalent distribution from the reconstructed 68Ga images while incorporating anatomical priors from the attenuation correction $\mu$-map. To assess correction accuracy, both global image comparisons and regional analyses were conducted to evaluate localized PR effects. The mean absolute error (MAE) and structural similarity index measure (SSIM) were used to quantify voxel-wise differences between corrected images and reference 18F PET images, ensuring that the PR correction restores radiotracer distribution while preserving structural integrity. All the results are given in Standardized Uptake Values relative (SUVr) to the average activity in the mouse.
@@ 2.6 Preclinical 68Ga PET studies
Experimental studies were conducted using the Inveon PET/CT system (Siemens Medical Solutions) at Memorial Sloan Kettering Cancer Center (New York, USA) to assess the feasibility of the proposed methods on the preclinical Inveon PET/CT scanner. The impact of the proposed PRCs were evaluated with two independent 68Ga acquisitions in mice. All PET images were reconstructed using GFIRST and corrected by random coincidences using a delayed time window approach within the acquisition [53], attenuation based on the values of the $\mu$-map and scatter using MCGPU-PET [54].
The first acquisition involved a preclinical AKP TV model to induce pulmonary tumor metastasis, as the lung is the first site of colonization when the injection is done into the tail vein [55]. To evaluate inflammation-associated tracer uptake, Feraheme (FH) nanoparticles were radiolabeled with 68Ga. FH is an iron oxide nanoparticle originally developed for the treatment of iron-deficiency anemia [56], but it also functions as a tracer in PET imaging due to its selective uptake by monocytes cancer or inflammation [57, 58]. The mouse received an intravenous injection of 22.755 MBq of 68Ga-FH, and PET imaging was initiated 3 hours post-injection, with a total scan duration of 30 minutes. The energy window for this acquisition was set to 350–700 keV. As respiratory gating was not applied, the resulting images may exhibit motion blur from lung movement, making tumor identification and quantification challenging.
The second set of experiments involved a preclinical prostate cancer xenograft model where PSMA-positive PC3-PIP cells were implanted subcutaneously into the right shoulder of immunocompromised mice and allowed to develop tumors for four weeks prior to PSMA-targeted PET imaging [59, 60, 61]. This model is widely used to study PSMA-specific therapies and diagnostics, as PC3-PIP cells express $\sim$2-fold higher PSMA levels compared to other engineered cell lines like LMD-PSMA [60]. The four-week growth period aligns with established protocols where tumor volumes typically reach measurable sizes [59]. The mouse was injected with 1.93 MBq of 68Ga-PSMA-617, and PET scanning began 8 minutes post-injection, lasting 50 minutes. The energy window for this acquisition was set to 350-650 keV.
@@ 3.1 Synthetic data analysis
[…]
@@ 4 Discussion
This study proposed and validated a DL-based PRC method using 3D RED-CNN architectures. The training framework incorporated PR in simulations with PenEasy [32] and MCGPU-PET [34] that were then reconstructed, ensuring the model learned PR effects under realistic imaging conditions. We proposed a novel loss function able to integrate the anatomical information based on the Mutual Information (MI) between the PET and the $\mu$-map. Furthermore, we evaluated the integration of the $\mu$-map as an extra channel within the training of the models. Two-Channel strategy introduced the $\mu$-map at the beginning of the training process. DualEncoder explores two parallel encoder chains for the PET and the $\mu$-map merging them before the decoder step. Finally, Single-Channel did not include the $\mu$-map in the training process and the information of the $\mu$-map was only introduced in the tissue-dependent loss function. According to the results for simulated studies presented in Fig 2 and Fig 3, the strategies including the $\mu$-map as an extra channel in the training outperformed the Single-Channel strategy, achieving an average SSIM 19.7% and 19.4% higher for Two-Channel and DualEnconder and respectively when compared to the PR corrupted input vs a 19.2% improvement in the Single-Channel. All the studied strategies led to enhanced results compared to RL deconvolution where the average CR enhancement was of only 9.9%. Moreover, as shown in Table 2, while the proposed methods based on DL successfully recovered the noise level of the target 18F images, RL deconvolution exhibited a noticeable increase (5.8%) in the noise. This observation is consistent with the behavior observed by standard deconvolution methods based on iterative algorithms [17].
@@
The incorporation of $\mu$-map anatomical information into the loss function enables the model to learn tissue-dependent PR effects. This approach allows for regionally adaptive adjustments, enhancing activity recovery in low-density regions while mitigating blurring in high-activity soft tissues. Integrating physics-informed neural network (PINN) techniques into the modified tissue-dependent loss function ensures that the correction process is guided by the underlying physical properties of PET imaging, rather than relying solely on data-driven learning [62, 63, 64]. By embedding tissue-dependent constraints, this approach improves quantitative accuracy, enhances structural consistency, and reduces the risk of artifacts, making it more robust to real PET acquisitions. These is particularly noticeable in the recovery of activity in low density regions such as lungs. Fig 4 shows how using the CNN-based approaches, CR in lungs was recovered with a variation 97% with respect to the ground truth for all of the studied networks, while conventional methods such as RL failed to recover the spillover in lungs with a variation of 77% with the ground truth, very similar to the PR corrupted images (83%). Also, while this study utilized $\mu$-maps from PET/CT, incorporating MRI-derived anatomical priors could further enhance PRC, particularly in PET/MR applications, where soft tissue contrast is superior. Future work should explore multi-modal strategies integrating MRI-based anatomical maps [65] to optimize PRC across different imaging modalities.
@@
Although MC–generated phantoms provide detailed anatomical and kinetic realism for training, they inevitably lack certain complexities of live small-animal PET data, such as respiratory and cardiac motion, detector non-uniformities, and unanticipated tracer heterogeneities. All CNN-based PRC methods outperformed RL-PRC when applied to simulated PET images, demonstrating superior contrast recovery, spatial preservation, and noise maintenance. However, their performance on real PET acquisitions faces the well-known challenge of generalizing models trained on synthetic data to real-world applications (simulation-to-reality gap) [66, 67]. The Two-Channel architecture retained crisp tissue boundaries and minimal artifacts, whereas the DualEncoder model showed degraded performance. We attribute these failures to two interrelated factors: (1) over-reliance on $\mu$-map features can suppress tracer uptake, leading to reduced functional contrast, and (2) the DualEncoder’s greater representational capacity and complex feature pathways can cause it to memorize simulation-specific noise characteristics, leading to diminished generalization in vivo. To mitigate these issues, future work will improve the current training set is to use real 18F PET/CT acquisitions (instead of numerical phantoms) as ground truth and simulate the corresponding 18F and 68Ga PET images under identical conditions to provide a more complete and realistic training dataset. Strategies such as domain adaptation, transfer learning, and hybrid training approaches that integrate real and synthetic data have been explored in other fields and could provide potential solutions [68, 69].
@@
In this work, the MI was included in the loss function, including a regularization parameter to determine the strength of the MI term in the loss function, since MI captures complex, non-linear relationships beyond simple intensity correlations. To determine the optimal MI weight $\lambda$, we performed an ablation study on our simulated validation set, evaluating $\lambda$ = 0 (no MI), 0.001, 0.01, and 0.1. We found that 0.01 provided the best compromise: lower values ($\lambda$ $<$ 0.01) failed to enforce sufficient anatomical consistency, while higher values ($\lambda$geq 0.1) resulted in “cross-talk” between PET and $\mu$-msp, manifesting as anatomical “hallucinations” that suppressed functional contrast and emphasized anatomical boundaries. By explicitly tuning $\lambda$, we ensure that the network leverages anatomical priors to guide the correction of positron-range blurring without overpowering the underlying radiotracer distribution.
@@
Although this study primarily focused on 3D RED-CNN architectures, the proposed tissue-dependent loss function has the potential to be integrated into other DL models, such as U-Net [29], Vision Transformers (ViT) [70, 71], and other modern NN architectures [72]. Future research will explore alternative methods for incorporating $\mu$-map information, including direct network integration, the use of attention mechanisms for selective feature enhancement [73], or multi-modal fusion techniques that combine PET and anatomical priors at different stages of the learning process [74]. A comparative analysis of these approaches could further optimize PRC by identifying the most effective way to balance functional and anatomical information.
@@ 5 Conclusion
  • Funding - We acknowledge support from the following grants: FASCINA (PID2021-126998OB-I00), Prototwin Project (TED2021-349130592B-I00), and 3PET (PDC2022-133057-I00), funded by the Spanish Ministry of Science, Innovation and Universities (MCIU) through AEI/10.13039/501100011033 and the European Union’s NextGenerationEU/PRT program. We also acknowledge funding from the TARTAGLIA Project (MIA.2021.M02.0005), supported by the Spanish Ministry of Economic Affairs and Digital Transformation as part of the Recovery, Resilience, and Transformation Plan, financed through the European Union’s NextGenerationEU funds. Additional support was provided by the National Institutes of Health (NIH), USA, under grants R01 EB03300, 1K99CA276804-01A1, and the Cancer Center Core Grant P30-CA008748. We also acknowledge funding from the TAU Project (PR47/21 TAU-CM), supported by the NextGenerationEU/PRT funds.

### v2

@@ 2.4 Implementation details
All experiments were conducted on an NVIDIA GeForce RTX 2080 Ti (12GB VRAM) using CUDA 12.2 and implemented in PyTorch neural network framework. The dataset was divided into 17 numerical phantoms for training, 8 for validation, and 20 for testing. Instead of using full-resolution images during training, a volumetric patch-based sampling strategy was adopted to optimize memory usage and enhance the model’s ability to capture localized PR effects. Each epoch included 4200 patches sampled for training and 1800 patches sampled for validation. Patches of three different sizes (24×24×24, 18×18×18, 12×12×12 voxels) were randomly extracted to promote multi-scale learning and improve generalization [moon2024m2former]. To prevent empty patches and improve robustness, a filtering step was applied to exclude empty regions based on the $\mu$-map. Given the physical constraints of PR effects, only random flipping in all three spatial directions (x, y, z) was applied. Scaling or zooming were avoided, as they could alter the PR effect, leading to inconsistencies in the network’s learning process.
Figure 2: Overview of the data preprocessing and sampling workflow. Each 3D image (18F, 68Ga, and $\mu$-map) was first clipped at the 99.99995th percentile to remove hot pixels. A normalization factor was then computed for each volume and stored without modifying the raw data. All datasets and their corresponding normalization factors were saved in an HDF5 file, grouped by numerical phantom (Digimouse or MOBY). During training, the DataLoader and custom Sampler dynamically read patches from the HDF5 file, apply the stored normalization factor, and perform random 3D flips before feeding them into the network.
@@
The network was optimized using the AdamW optimizer [loshchilov2017decoupled], with an initial learning rate of $10^{-3}$, and a OneCycleLR scheduler [smith2019super] dynamically adjusting learning rates and momentum to enhance training stability. The batch size was set to 120. Early stopping was applied with a patience of 50 epochs and a minimum required improvement of $10^{-7}$ in the validation loss.
@@
To handle large volumetric PET datasets efficiently, the HDF5 file format [neonhdf5] was used for data storage and sampling, allowing efficient random patch extraction without loading all the patches into memory, faster data access, and reduced I/O bottlenecks during training, as well as optimized GPU memory usage [bieder2023diffusion], enabling training on hardware with limited VRAM. Unlike some implementations that use mixed precision training to reduce memory usage, our approach was trained using full precision (FP32) to ensure numerical stability, preventing potential precision errors.
For each PET volume, voxel intensities above the 99.99995th percentile were clipped to avoid hot pixels due to reconstruction. Each PET and u-map image was then normalized by its mean intensity, yielding mean-equalized volumes with comparable intensity ranges across channels. The scaling coefficients were stored for later use during the training (see Figure 2)
@@
In this work, we leverage the properties of convolutional layers and their receptive fields to infer full images without patching [luo2016understanding]. Convolutional operations learn spatial patterns locally and apply them uniformly across larger images due to shared weights and growing receptive fields. Additionally, spatial equivariance ensures that features learned in one region remain valid elsewhere [cohen2019general]. By integrating volumetric patches during training and inferring full images during testing, our approach enhances efficiency while maintaining consistent feature extraction across the entire volume.
@@ 2.5 Validation
Validation was performed using simulated PET datasets, where both 68Ga and 18F PET images were generated under same conditions. The network was trained to predict the reconstructed 18F-equivalent distribution from the reconstructed 68Ga images while incorporating anatomical priors from the $\mu$-map. To assess correction accuracy, both global image comparisons and regional analyses were conducted to evaluate localized PR effects. The mean absolute error (MAE) and structural similarity index measure (SSIM) were used to quantify voxel-wise differences between corrected images and reference 18F PET images, ensuring that the PR correction restores radiotracer distribution while preserving structural integrity. All the results are given in Standardized Uptake Values relative (SUVr) to the average activity in the mouse.
@@ 2.6 Preclinical 68Ga PET studies
Experimental studies were conducted using the Inveon PET/CT system (Siemens Medical Solutions) at Memorial Sloan Kettering Cancer Center (New York, USA) to assess the feasibility of the proposed methods on the preclinical Inveon PET/CT scanner. The impact of the proposed PRCs were evaluated with two independent 68Ga acquisitions in mice. All PET images were reconstructed using GFIRST and corrected by random coincidences using a delayed time window approach within the acquisition [brasse2005correction], attenuation based on the values of the $\mu$-map and scatter using MCGPU-PET [cabello2025comparison].
The first acquisition involved a preclinical AKP TV model to induce pulmonary tumor metastasis, as the lung is the first site of colonization when the injection is done into the tail vein [thies2020pathological]. To evaluate inflammation-associated tracer uptake, Feraheme (FH) nanoparticles were radiolabeled with 68Ga. FH is an iron oxide nanoparticle originally developed for the treatment of iron-deficiency anemia [coyne2009ferumoxytol], but it also functions as a tracer in PET imaging due to its selective uptake by monocytes cancer or inflammation [yuan2018heat, gholami2020radio]. The mouse received an intravenous injection of 22.755 MBq of 68Ga-FH, and PET imaging was initiated 3 hours post-injection, with a total scan duration of 30 minutes. The energy window for this acquisition was set to 350–700 keV. As respiratory gating was not applied, the resulting images may exhibit motion blur from lung movement, making tumor identification and quantification challenging.
The second set of experiments involved a preclinical prostate cancer xenograft model where PSMA-positive PC3-PIP cells were implanted subcutaneously into the right shoulder of immunocompromised mice and allowed to develop tumors for four weeks prior to PSMA-targeted PET imaging [boinapally2021prostate, mukherjee2016development, dam2017psma]. This model is widely used to study PSMA-specific therapies and diagnostics, as PC3-PIP cells express $\sim$2-fold higher PSMA levels compared to other engineered cell lines like LMD-PSMA [mukherjee2016development]. The four-week growth period aligns with established protocols where tumor volumes typically reach measurable sizes [boinapally2021prostate]. The mouse was injected with 1.93 MBq of 68Ga-PSMA-617, and PET scanning began 8 minutes post-injection, lasting 50 minutes. The energy window for this acquisition was set to 350-650 keV.
@@ 3.1 Synthetic data analysis
Figure 5: A CR ($\%$) and B CNR comparison across organs -myocardium, bladder, lungs and liver- and methods. Abbreviations: RL-PRC=Richardson-Lucy Positron Range Correction, CR=Contrast Recovery, CNR=Contrast-to-Noise Ratio,
@@ 4 Discussion
This study proposed and validated a DL-based PRC method using 3D RED-CNN architectures. The training framework incorporated PR in simulations with PenEasy [sempau2006configuration] and MCGPU-PET [herraiz2024mcgpu] that were then reconstructed, ensuring the model learned PR effects under realistic imaging conditions. We proposed a novel loss function able to integrate the anatomical information based on the Mutual Information (MI) between the PET and the $\mu$-map. Furthermore, we evaluated the integration of the $\mu$-map as an extra channel within the training of the models. Two-Channel strategy introduced the $\mu$-map at the beginning of the training process. DualEncoder explores two parallel encoder chains for the PET and the $\mu$-map merging them before the decoder step. Finally, Single-Channel did not include the $\mu$-map in the training process and the information of the $\mu$-map was only introduced in the tissue-dependent loss function. According to the results for simulated studies presented in Fig 3 and Fig 4, the strategies including the $\mu$-map as an extra channel in the training outperformed the Single-Channel strategy, achieving an average SSIM 19.7% and 19.4% higher for Two-Channel and DualEnconder and respectively when compared to the PR corrupted input vs a 19.2% improvement in the Single-Channel. All the studied strategies led to enhanced results compared to RL deconvolution where the average CR enhancement was of only 9.9%. Moreover, as shown in Table 2, while the proposed methods based on DL successfully recovered the noise level of the target 18F images, RL deconvolution exhibited a noticeable increase (5.8%) in the noise. This observation is consistent with the behavior observed by standard deconvolution methods based on iterative algorithms [gavriilidis2022positron].
@@
The incorporation of $\mu$-map anatomical information into the loss function enables the model to learn tissue-dependent PR effects. This approach allows for regionally adaptive adjustments, enhancing activity recovery in low-density regions while mitigating blurring in high-activity soft tissues. Integrating physics-informed neural network (PINN) techniques into the modified tissue-dependent loss function ensures that the correction process is guided by the underlying physical properties of PET imaging, rather than relying solely on data-driven learning [raissi2019physics, karniadakis2021physics, weikun2023physics]. By embedding tissue-dependent constraints, this approach improves quantitative accuracy, enhances structural consistency, and reduces the risk of artifacts, making it more robust to real PET acquisitions. These is particularly noticeable in the recovery of activity in low density regions such as lungs. Fig 5 shows how using the CNN-based approaches, CR in lungs was recovered with a variation 97% with respect to the ground truth for all of the studied networks, while conventional methods such as RL failed to recover the spillover in lungs with a variation of 77% with the ground truth, very similar to the PR corrupted images (83%). Also, while this study utilized $\mu$-maps from PET/CT, incorporating MRI-derived anatomical priors could further enhance PRC, particularly in PET/MR applications, where soft tissue contrast is superior. Future work should explore multi-modal strategies integrating MRI-based anatomical maps [vandenberghe2015pet] to optimize PRC across different imaging modalities.
Both Digimouse and MOBY phantoms were used for training and testing to promote domain-invariant learning of PRC. Although both represent murine anatomy, they differ in structural detail and $\mu$-map origin: Digimouse $\mu$-maps are derived from real CT scans of an ex-vivo mouse, whereas MOBY $\mu$-maps are generated from an anatomical atlas with smoother, idealized tissue boundaries. To ensure compatibility, both datasets were harmonized to match the voxel size and simulation parameters of the Siemens Inveon PET system. This harmonization, while necessary, may have slightly reduced the structural detail of Digimouse and contributed to the minor residual differences observed, since more Digimouse cases were used for training, while only MOBY was used for evaluation. Nevertheless, the quantitative results (RMSE and SSIM) confirmed that the proposed method remains robust and generalizes well across domains.
The normalization strategy also plays an important role in ensuring stable and comparable learning across network architectures. The Single-Channel network, composed only of convolutional and ReLU layers, is inherently scale-independent. This property arises from the shift-invariant nature of convolutions [lecun2015deep] and the positive homogeneity of ReLU activations, which allow the model to preserve proportional relationships in the input without requiring explicit normalization during inference. In contrast, the Two-Channel and DualEncoder architectures jointly process PET and $\mu$-map inputs, which differ in intensity distribution. Without normalization, these disparities could introduce bias in feature learning and affect convergence stability. By normalizing each input by its mean intensity, both modalities are brought to comparable scales, ensuring balanced contribution during training and consistent feature representation across channels.
@@
Although MC–generated phantoms provide detailed anatomical and kinetic realism for training, they inevitably lack certain complexities of live small-animal PET data, such as respiratory and cardiac motion, detector non-uniformities, and unanticipated tracer heterogeneities. All CNN-based PRC methods outperformed RL-PRC when applied to simulated PET images, demonstrating superior contrast recovery, spatial preservation, and noise maintenance. However, their performance on real PET acquisitions faces the well-known challenge of generalizing models trained on synthetic data to real-world applications (simulation-to-reality gap) [tremblay2018training, alkhalifah2022mlreal]. The Two-Channel architecture retained crisp tissue boundaries and minimal artifacts, whereas the DualEncoder model showed degraded performance. We attribute these failures to two interrelated factors: (1) over-reliance on $\mu$-map features can suppress tracer uptake, leading to reduced functional contrast, and (2) the DualEncoder’s greater representational capacity and complex feature pathways can cause it to memorize simulation-specific noise characteristics, leading to diminished generalization in vivo. To mitigate these issues, future work will improve the current training set is to use real 18F PET/CT acquisitions (instead of numerical phantoms) as ground truth and simulate the corresponding 18F and 68Ga PET images under identical conditions to provide a more complete and realistic training dataset. Strategies such as domain adaptation, transfer learning, and hybrid training approaches that integrate real and synthetic data have been explored in other fields and could provide potential solutions [tobin2017domain, zhao2020sim].
For real 68Ga-FH and 68Ga-PSMA-617 acquisitions (Figure 6), accuracy was ensured by preserving the total number of detected counts, so that activity was only spatially redistributed. Compared with the Richardson–Lucy correction, the proposed method recovered physiologically plausible uptake in the tumor region without introducing spurious activity. Although ideal validation would involve paired acquisitions with isotopes with large and small PR, this is experimentally challenging because differences in tracer biodistribution may occur, even when the same tracer is used. Expert review by nuclear medicine physicians at MSKCC confirmed that the corrected images exhibited realistic uptake patterns, supporting the quantitative reliability of the proposed approach.
@@
In this work, the MI was included in the loss function, including a regularization parameter to determine the strength of the MI term in the loss function, since MI captures complex, non-linear relationships beyond simple intensity correlations. To determine the optimal MI weight $\lambda$, we performed an ablation study on our simulated validation set, evaluating $\lambda$ = 0 (no MI), 0.001, 0.01, and 0.1. We found that 0.01 provided the best compromise: lower values ($\lambda$ << 0.01) failed to enforce sufficient anatomical consistency, while higher values ($\lambda$ $\geq$ 0.1) resulted in “cross-talk” between PET and $\mu$-msp, manifesting as anatomical “hallucinations” that suppressed functional contrast and emphasized anatomical boundaries. By explicitly tuning $\lambda$, we ensure that the network leverages anatomical priors to guide the correction of positron-range blurring without overpowering the underlying radiotracer distribution.
@@
Although this study primarily focused on 3D RED-CNN architectures, the proposed tissue-dependent loss function has the potential to be integrated into other DL models, such as U-Net [ronneberger2015u], Vision Transformers (ViT) [dosovitskiy2020image, takahashi2024comparison], and other modern NN architectures [abdou2022literature]. Future research will explore alternative methods for incorporating $\mu$-map information, including direct network integration, the use of attention mechanisms for selective feature enhancement [vaswani2017attention], or multi-modal fusion techniques that combine PET and anatomical priors at different stages of the learning process [huang2020review]. A comparative analysis of these approaches could further optimize PRC by identifying the most effective way to balance functional and anatomical information.
@@ 5 Conclusion
  • Funding - We acknowledge support from the following grants: FASCINA (PID2021-126998OB-I00), Prototwin Project (TED2021-349130592B-I00), and 3PET (PDC2022-133057-I00), funded by the Spanish Ministry of Science, Innovation and Universities (MCIU) through AEI/10.13039/501100011033 and the European Union’s NextGenerationEU/PRT program. This work received support within the framework of the LUNABRAIN-CM (TEC-2024/TEC-43) project, funded by Comunidad de Madrid (Spain) through the R&\&D activities program in technologies, granted by Order 5696/2024. We also acknowledge funding from the TARTAGLIA Project (MIA.2021.M02.0005), supported by the Spanish Ministry of Economic Affairs and Digital Transformation as part of the Recovery, Resilience, and Transformation Plan, financed through the European Union’s NextGenerationEU funds. Additional support was provided by the National Institutes of Health (NIH), USA, under grants R01EB033000, 1K99CA276804-01A1, and the Cancer Center Core Grant P30-CA008748. We also acknowledge funding from the TAU Project (PR47/21 TAU-CM), supported by the NextGenerationEU/PRTR funds.